# Large Language Models are Easily Confused: A Quantitative Metric, Security Implications and Typological Analysis

### v1

@@ Abstract
Language Confusion is a phenomenon where Large Language Models (LLMs) generate text that is neither in the desired language, nor in a contextually appropriate language. This phenomenon presents a critical challenge in text generation by LLMs, often appearing as erratic and unpredictable behavior. We hypothesize that there are linguistic regularities to this inherent vulnerability in LLMs and shed light on patterns of language confusion across LLMs. We introduce a novel metric, Language Confusion Entropy, designed to directly measure and quantify this confusion, based on language distributions informed by linguistic typology and lexical variation. Comprehensive comparisons with the Language Confusion Benchmark  (Marchisio et al., 2024) confirm the effectiveness of our metric, revealing patterns of language confusion across LLMs. We further link language confusion to LLM security, and find patterns in the case of multilingual embedding inversion attacks. Our analysis demonstrates that linguistic typology offers theoretically grounded interpretation, and valuable insights into leveraging language similarities as a prior for LLM alignment and security  [^1] The language graphs for language similarities and code are publicly available https://github.com/siebeniris/QuantifyingLanguageConfusion/. .
@@ A Quantitative Metric, Security Implications and Typological Analysis
Yiyi Chen1, QiongXiu Li2, Russa Biswas1, Johannes Bjerva1, 1Department of Computer Science 2Department of Electronic Systems Aalborg University, Copenhagen, Denmark {yiyic,rubi,jbjerva}@cs.aau.dk, qili@es.aau.dk
@@ 1 Introduction
Multilingual Large Language Models (LLMs) revolutionized Natural Language Processing (NLP), offering crosslinguality in various applications, including translation Zhu et al. (2024), text generation Chen et al. (2022), and information retrieval Guo et al. (2024). Besides the challenges faced by LLMs such as bias and fairness Talat et al. (2022), hallucinations Augenstein et al. (2024), multilingual LLMs are more vulnerable to adversarial and inversion attacks than monolingual LLMs Song et al. (2024); Chen et al. (2024a, b).
Multilingual LLMs are trained on a diverse range of linguistic data to represent the intricacies of multiple languages within a single model. However, this often results in inconsistencies in comprehension and response, leading to language confusion – instances where LLMs generate text that is neither in the desired language nor in a contextually appropriate one. For example, when an LLM is queried/prompted in Arabic, it may respond in text that is either partially or entirely in languages other than Arabic, e.g., English.
Marchisio et al. (2024) propose metrics to measure the percentage of model responses containing no undesired languages at both line and word levels but fail to capture nuances within language distributions. It is observed that language confusion tends to occur when the model’s distribution over the next tokens is flat. We hypothesize that language confusion in LLMs is not merely a performance limitation but an inherent vulnerability, partly due to imbalanced multilingual data sources, which are amplified with increasing numbers of languages and can be analyzed through linguistic typology.
@@
RQ2: How do linguistic similarities influence language confusion, and how can this knowledge be applied to enhance LLM alignment and security?
To this end, we propose a novel metric called Language Confusion Entropy, which provides a quantifiable measure of uncertainty and facilitates the detection of when an LLM is confused. Building on observations by Marchisio et al. (2024) that uniformity of the distribution indicates higher uncertainty, Language Confusion Entropy re-weights language distributions by emphasizing long-tail distributions, effectively capturing language confusion in multilingual LLM generation tasks. Furthermore, we demonstrate that this metric uncovers patterns of language confusion during both the training and evaluation phases of multilingual inversion attacks (Chen et al., 2024a).
In addition, we construct language graphs based on linguistic typology to analyze language confusion, revealing a strong correlation between language confusion and semantic similarities between languages. Our analysis shows that low-resource languages exhibit less confusion, while training across diverse scripts and language families mitigates language confusion more effectively than training within the same script or language family in inversion attacks. These findings indicate that leveraging language similarities grounded in linguistic typology could serve as a valuable prior for enhancing LLM alignment and security. Our main contributions are as follows:
1) We propose a novel metric Language Confusion Entropy to measure language confusion in LLMs considering the nuances of language distributions. To the best of our knowledge, we are the first to quantify language confusion.
@@ 2 Related Works
The phenomenon of language confusion has not only been pervasive in LLMs, but also in tasks pertinent to LLM security, such as multilingual inversion attacks (Chen et al., 2024b, a). Furthermore, Chen et al. (2024a) observes language confusion across 20 languages from diverse scripts and language families in multilingual embedding inversion. They analyzed the pattern of confusion using basic typological features between train and eval languages, however, the proposed Language Confusion Entropy provides more interpretable analysis.
@@
Backdoor attacks on multilingual machine translation pose significant threats, as injecting very less poisoned data into low-resource language pairs can achieve a high attack success rate (ASR) in high-resource language pairs (Wang et al., 2024). Poisoning instruction-tuning data for one or two languages can affect other languages, surpassing 99% ASR in the cross-lingual settings in prominent LLMs resisting current defenses (He et al., 2024).
Multilingual textual embedding inversion attacks pose additional risks, as any encoder can be attacked to reconstruct original texts. Traditional defenses for monolingual LLMs are ineffective for multilingual LLMs (Chen et al., 2024b, a). Moreover, Song et al. (2024) generates language blending for adversarial attacks, necessitating systematic analysis of language similarity and language confusion for targeted defenses.
@@ 3 Explainable Language Confusion
To investigate the phenomenon of language confusion, we use the datasets i) Language Confusion Benchmark (LCB) (Marchisio et al., 2024) and ii) Multilingual Textual Embedding Inversion (MTEI) (Chen et al., 2024a) (see Appendix for task details and Table 8 for processed datasets sample).
@@
Generation Settings Consider a LLM trained or prompted with a set of $n\in\mathbf{N}$ languages $L_{s}=\{l_{1},\cdots,l_{n}\}$, and $l_{t}$ is the target language. We probe language confusion for both LLM instruction and textual embedding inversion attacks
Monolingual Generation LCB: The model is queried in language $l_{t}$, and the response is expected in $l_{t}$. MTEI: The inversion model, trained on $l_{t}$, inverts embeddings in $l_{t}$. Here, $l_{t}\in L_{s}$, i.e., the evaluated language is part of the training languages.
Crosslingual Generation LCB: The model is instructed in language $l_{s}$ to provide a response in $l_{t}$, where $l_{t}\neq l_{s}$. MTEI: The inversion model, trained on $L_{s}$, inverts embeddings in $l_{t}$. In this setting, $l_{t}\notin L_{s}$, meaning the evaluated language differs from the training languages.
Quantifying Language Confusion
@@
Language Identification
Language confusion can occur at both the word level and the line level. To measure it, we use existing language identification (LID) tools. Initially, we employ Lingua to detect languages as it offers the highest accuracy among existing LID tools, particularly for word-level detection. [^2] https://github.com/pemistahl/lingua-py However, since Lingua supports only 75 languages, we supplement it with fastText (Joulin et al., 2016) (which supports 176 languages) for languages unidentified by Lingua.
Line-level Detection We split the generated output into lines by newline characters and detect the language of each line.
Word-level Detection To detect languages more accurately at the word level, we tokenize the text at the line level using language-specific tokenizers such as jieba (Sun et al., 2013) (Chinese), Hebrew Tokenizer (Levin and Oriyan, 2018) Kiwipiepy (Lee, 2024) (Korean), fugashi (McCann, 2020) (Japanese), and NLTK Word Tokenizer  (Bird and Loper, 2004) for other languages. We then identify the language of each tokenized word.
Finally, we compile the detected languages into distributions at both word and line levels for further analysis. Examples of the pre-processed data are presented in Table 8. Language confusion matrices are then constructed by applying language confusion entropy to these distributions (ref. Fig. 1  1).
@@
Large Language Models
The evaluated LLMs in LCB include Command R (35B parameters) Command R+ (104B) GPT-3.5 Turbo and GPT-4 Turbo Mistral Large Mistral 8x7B LLaMA 2 70B Instruct and LLaMA 3 70B Instruct while in  MTEI the inversion models are trained with mT5 (580M) (Xue, 2020) and multilingual-e5-base (me5) (580M). (See Table 5 for details of LLMs.)
Language Graphs from Linguistic Typology
We construct language graphs from a diverse range of typological features, such as colexification patterns (Rzymski et al., 2020; Fellbaum, 2010; Navigli and Ponzetto, 2010), lexicon (Wichmann et al., 2012), pathological and morphological-syntactical features (Haspelmath, 2008; Skirgård et al., 2023), as well as from a collection of existing language embeddings trained from NLP tasks incorporating linguistic typology (Östling and Tiedemann, 2017; Östling and Kurfalı, 2023; Chen et al., 2023b). We then generate language similarity matrices from the language graphs by calculating pairwise similarity using either Jaccard Index or Cosine Similarity (ref. Fig. 1  2). (See details in Appendix B).
The Role of Language Similarity in Language Confusion
@@
Table 1: Spearman correlation between Language Confusion Entropy and Pass Rates at both Word level and Line level with LCB. Strongest correlation is in bold.
4 Analysis and Results
4.1 Language Confusion in LLM Prompting
Language Confusion Entropy vs. Pass Rates
Compared to Marchisio et al. (2024), our approach detects language confusion across all languages, including at the word level, by using language-specific tokenizers and more accurate language identification (LID) tool. We reproduce LPR and WPR (ref. Table 1011) and compute $H_{\mathbf{C}}$[L] and $H_{\mathbf{C}}$[W] (ref. Table 9) in both crosslingual and monolingual settings, for 14 languages and 8 LLMs, following (Marchisio et al., 2024).
To evaluate the efficacy of language confusion entropy compared to pass rate metrics, we calculate the Spearman correlation [^3] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.spearmanr.html coefficients between these metrics across levels and generation settings. Overall, $H_{\mathbf{C}}$[L] shows a strong negative correlation with LPR across all generation settings, On the other hand, $H_{\mathbf{C}}$[W] - which is based on more detailed language distributions - exhibits a weaker correlation with WPR, as WPR only considers English words in non-Latin script languages. Despite the weaker correlation, it remains statistically significant for all languages, especially crosslingually.
@@ 4.1 Language Confusion in LLM Prompting
Observing language confusion across datasets at the line level for crosslingual settings (Fig. $4$We use “Language Confusion Entropy” and “Confusion Entropy” interchangeably in this paper.), a clear pattern emerges. Complex Prompts has the highest confusion across all languages, while ShareGPT shows the lowest confusion for most languages, except for French and Spanish. Consistent with previous findings, languages written in non-Latin scripts show higher confusion in datasets like Okapi and ShareGPT. However, in Complex Prompts, non-Latin-script languages such as Chinese, Korean, Arabic, and Japanese demonstrate lower confusion than Latin-script languages.
@@ 4.2 Language Confusion in Multilingual Textual Embedding Inversion Security
Moreover, $H_{\mathbf{C}}$[L] and $H_{\mathbf{C}}$[W] are both strongly correlated with the proportion of languages in pre-training data of mT5, particularly in crosslingual setting. This indicates that languages with higher representation in the pre-training data are more prone to confusion, both at the word and line levels. This observation is further validated by empirical evidence: when an inversion model trained on Arabic is used to invert embeddings in Meadow Mari (unseen in mT5), the model often generates English, highlighting the influence of pre-training data on embedding inversion attacks. Fig 6 (top) provides a visualization of language confusion at the line level for each language, alongside their corresponding reconstruction performance at each step of the evaluation in the crosslingual inversion attacks. It shows directly that lower-resourced languages present lower confusion, especially for non-Latin script languages, whereas they are also more vulnerable in terms of higher reconstruction performance, for example, Gujarati, Punjabi, and Urdu.
Language Confusion Entropy for Train Languages
Our study reveals that inversion performance significantly improves when trained in in-script/in-family settings(ref. Table 13 in Appendix). Crosslingual inversion performances are comparable to in-script/in-family training when trained in Kazakh (Latin-script) combined with Gujarati and Punjabi, respectively, and language confusion is notably lower. This suggests that while similar languages tend to increase confusion, certain crosslingual combinations can achieve strong performance without the added confusion seen in in-family/in-script training. Overall, these findings highlight the trade-off between inversion performance and language confusion indicating further optimization is needed to strike the ideal balance between them.
@@ 4.3 Language Confusion and Linguistic Typology
Table 3 shows the best results from KL divergence between language confusion and language similarities based on different language graphs for both LCB and MTEI, using Algorithm 1, the whole results are presented in 15 in Appendix.
@@
4.4 Discussion

### v2

@@ Abstract
Language Confusion is a phenomenon where Large Language Models (LLMs) generate text that is neither in the desired language, nor in a contextually appropriate language. This phenomenon presents a critical challenge in text generation by LLMs, often appearing as erratic and unpredictable behavior. We hypothesize that there are linguistic regularities to this inherent vulnerability in LLMs and shed light on patterns of language confusion across LLMs. We introduce a novel metric, Language Confusion Entropy, designed to directly measure and quantify this confusion, based on language distributions informed by linguistic typology and lexical variation. Comprehensive comparisons with the Language Confusion Benchmark  (Marchisio et al., 2024) confirm the effectiveness of our metric, revealing patterns of language confusion across LLMs. We further link language confusion to LLM security and find patterns in the case of multilingual embedding inversion attacks. Our analysis demonstrates that linguistic typology offers theoretically grounded interpretation, and valuable insights into leveraging language similarities as a prior for LLM alignment and security. [^1] The language graphs for language similarities and code are publicly available https://github.com/siebeniris/QuantifyingLanguageConfusion/
@@ A Quantitative Metric, Security Implications and Typological Analysis
Yiyi Chen1, QiongXiu Li2, Russa Biswas1, Johannes Bjerva1 1Department of Computer Science, 2Department of Electronic Systems Aalborg University, Copenhagen, Denmark {yiyic, rubi, jbjerva}@cs.aau.dk, qili@es.aau.dk
@@ 1 Introduction
Multilingual Large Language Models (LLMs) revolutionized Natural Language Processing (NLP), offering crosslinguality in various applications, including translation Zhu et al. (2024), text generation Chen et al. (2022), and information retrieval Guo et al. (2024). Besides the challenges faced by LLMs such as bias and fairness Talat et al. (2022), hallucinations Augenstein et al. (2024), multilingual LLMs are more vulnerable to adversarial and inversion attacks than monolingual LLMs Song et al. (2024); Chen et al. (2025, 2024).
Multilingual LLMs are trained on data in a diverse range of languages to represent the intricacies of multiple languages within a single model. However, this often results in inconsistencies in comprehension and response, leading to language confusion – instances where LLMs generate text that is neither in the desired language nor in a contextually appropriate one. For example, when an LLM is queried/prompted in Arabic, it may respond in text that is either partially or entirely in languages other than Arabic, e.g., English.
Marchisio et al. (2024) propose metrics to measure the percentage of model responses containing no undesired languages at both line and word levels but fail to capture nuances within language distributions. It is observed that language confusion tends to occur when the model’s distribution over the next tokens is flat. We hypothesize that language confusion in LLMs is not merely a performance limitation but an inherent vulnerability, partly due to imbalanced pre-training multilingual data sources, which are amplified with increasing numbers of languages and can be analyzed through language similarities derived from linguistic typology and other resources.
@@
RQ2: How do language similarities influence language confusion, and how can this knowledge be applied to enhance LLM alignment and security?
To this end, we propose a novel metric called Language Confusion Entropy, which provides a quantifiable measure of uncertainty and facilitates the detection of when an LLM is confused. Building on observations by Marchisio et al. (2024) that uniformity of the distribution indicates higher uncertainty, Language Confusion Entropy re-weights language distributions by emphasizing long-tail distributions, effectively capturing language confusion in multilingual LLM generation tasks. Furthermore, we demonstrate that this metric uncovers patterns of language confusion during both the training and evaluation phases of multilingual inversion attacks (Chen et al., 2025).
In addition, we construct language graphs based on language similarities derived from linguistic resources to analyze language confusion, revealing a strong correlation between language confusion and semantic similarities between languages. Our analysis shows that low-resource languages exhibit less confusion while training across diverse scripts and language families mitigates language confusion more effectively than training within the same script or language family in inversion attacks. These findings indicate that leveraging language similarities grounded in linguistic resources could serve as a valuable prior for enhancing LLM alignment and security. Our main contributions are as follows:
1) We propose a novel metric Language Confusion Entropy to measure language confusion in LLMs considering the nuances of language distributions. To the best of our knowledge, we are the first to quantify language confusion probabilistically.
@@ 2 Related Works
The phenomenon of language confusion has not only been pervasive in LLMs, but also in tasks pertinent to LLM security, such as multilingual inversion attacks (Chen et al., 2024, 2025). Furthermore, Chen et al. (2025) observes language confusion across 20 languages from diverse scripts and language families in multilingual embedding inversion. They analyzed the pattern of confusion using basic typological features between train and eval languages with regression analysis, in comparison, the proposed Language Confusion Entropy provides a more interpretable analysis.
@@
Backdoor attacks on multilingual machine translation pose significant threats, as injecting poisoned data into low-resource language pairs can achieve a high attack success rate (ASR) in high-resource language pairs (Wang et al., 2024). Poisoning instruction-tuning data for one or two languages can affect other languages, surpassing 99% ASR in the cross-lingual setting in prominent LLMs resisting current defenses (He et al., 2024).
Multilingual textual embedding inversion attacks pose additional risks, as any encoder can be attacked to reconstruct original texts. Traditional defenses for monolingual LLMs are ineffective for multilingual LLMs (Chen et al., 2024, 2025). Moreover, Song et al. (2024) generates language blending for adversarial attacks, necessitating systematic analysis of language similarity and language confusion for targeted defenses.
@@ 3 Explainable Language Confusion
To investigate the phenomenon of language confusion, we use the datasets i) Language Confusion Benchmark (LCB) (Marchisio et al., 2024) and ii) Multilingual Textual Embedding Inversion (MTEI) (Chen et al., 2025) (see Appendix for task details and Table 8 for processed datasets sample).
@@
3.1 Generation Settings
Consider a LLM trained or prompted with a set of $n\in\mathbb{N}$ source languages $L_{s}=\{l_{1},\cdots,l_{n}\}$, and $l_{t}$ is the target language. We probe language confusion for both LLM instruction and textual embedding inversion attacks.
Monolingual Generation
LCB: The model is queried in language $l_{t}$, and the response is expected in $l_{t}$. MTEI: The inversion model, trained on $l_{t}$, inverts embeddings in $l_{t}$. Here, $l_{t}\in L_{s}$, i.e., the evaluated language is part of the training languages.
Crosslingual Generation
LCB: The model is instructed in language $l_{s}$ to provide a response in $l_{t}$, where $l_{t}\neq l_{s}$. MTEI: The inversion model, trained on $L_{s}$, inverts embeddings in $l_{t}$. In this setting, $l_{t}\notin L_{s}$, meaning the evaluated language differs from the training languages.
3.2 Quantifying Language Confusion
@@
3.3 Language Identification
Language confusion can occur at both the word level and the line level. To measure it, we use existing language identification (LID) tools. Initially, we employ Lingua to detect languages as it offers the highest accuracy among existing LID tools, particularly at word level. [^2] https://github.com/pemistahl/lingua-py However, since Lingua supports only 75 languages, we supplement it with fastText (Joulin et al., 2016) (which supports 176 languages) for languages unidentified by Lingua.
Line-level Detection
We split the generated output into lines by newline characters and detect the language of each line.
Word-level Detection
To detect languages more accurately at the word level, we first tokenize the text at the line level using language-specific tokenizers such as jieba (Sun et al., 2013) (Chinese), Hebrew Tokenizer (Levin and Oriyan, 2018) Kiwipiepy (Lee, 2024) (Korean), fugashi (McCann, 2020) (Japanese), and NLTK Word Tokenizer  (Bird and Loper, 2004) for other languages. We then identify the language of each tokenized word.
Finally, we compile the detected languages into distributions at both word and line levels for further analysis. Examples of the pre-processed data are presented in Table 8. Language confusion matrices are then constructed by applying language confusion entropy to these distributions, and the results are aggregated per language (ref. Fig. 1  1).
@@ 3.3 Language Identification
3.4 Large Language Models
The evaluated LLMs in LCB include Command R (35B parameters),Command R+ (104B), GPT-3.5 Turbo (Brown, 2020), and GPT-4 Turbo (Achiam et al., 2023), Mistral Large, Mistral 8x7B (Jiang et al., 2024), LLaMA 2 70B Instruct (Touvron et al., 2023), and LLaMA 3 70B Instruct, while in  MTEI the inversion models are trained with mT5 (580M) (Xue, 2020) and multilingual-e5-base (me5) (580M). (See Table 5 for details of LLMs).
3.5 Language Graphs
We construct language graphs from a diverse range of typological features, such as colexification patterns (Rzymski et al., 2020; Fellbaum, 2010; Navigli and Ponzetto, 2010), lexicon (Wichmann et al., 2012), phonological and morphological-syntactical features (Haspelmath, 2008; Skirgård et al., 2023), as well as from a collection of existing language embeddings trained from NLP tasks incorporating linguistic typology (Östling and Tiedemann, 2017; Östling and Kurfalı, 2023; Chen et al., 2023b). We then generate language similarity matrices from the language graphs by calculating pairwise similarity using either Jaccard Index or Cosine Similarity (ref. Fig. 1  2). (See details in Appendix B).
3.6 The Role of Language Similarity in Language Confusion
@@
4 Analysis and Results
4.1 Language Confusion in LLM Prompting
Language Confusion Entropy vs. Pass Rates
Compared to Marchisio et al. (2024), our approach detects language confusion across all languages, including at the word level, by using language-specific tokenizers and a more accurate language identification (LID) tool. We reproduce LPR and WPR (ref. Table 1011) and compute $H_{\mathbf{C}}$[L] and $H_{\mathbf{C}}$[W] (ref. Table 9) in both crosslingual and monolingual settings, for 14 languages and 8 LLMs, following (Marchisio et al., 2024).
To evaluate the efficacy of language confusion entropy compared to pass rate metrics, we calculate the Spearman correlation [^3] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.spearmanr.html coefficients between these metrics across levels and generation settings. Overall, $H_{\mathbf{C}}$[L] shows a strong negative correlation with LPR across all generation settings. Moreover, $H_{\mathbf{C}}$[W] - which is based on more detailed language distributions - exhibits a weaker correlation with WPR, as WPR only considers English words in non-Latin script languages. Despite the weaker correlation, it remains statistically significant for all languages, especially crosslingually.
@@ 4.1 Language Confusion in LLM Prompting
Table 1: Spearman correlation between Language Confusion Entropy and Pass Rates at both Word level and Line level with LCB. The strongest correlation is in bold.
@@
Observing language confusion across datasets at the line level for crosslingual settings (Fig. 5), [^4] We use “Language Confusion Entropy” and “Confusion Entropy” interchangeably in this paper. a clear pattern emerges. Complex Prompts has the highest confusion across all languages, while ShareGPT shows the lowest confusion for most languages, except for French and Spanish. Consistent with previous findings, languages written in non-Latin scripts show higher confusion in datasets like Okapi and ShareGPT. However, in Complex Prompts, non-Latin-script languages such as Chinese, Korean, Arabic, and Japanese demonstrate lower confusion than Latin-script languages.
@@ 4.2 Language Confusion in Multilingual Textual Embedding Inversion Security
Moreover, $H_{\mathbf{C}}$[L] and $H_{\mathbf{C}}$[W] are both strongly correlated with the proportion of languages in pre-training data of mT5, particularly in crosslingual setting. This indicates that languages with higher representation in the pre-training data are more prone to confusion, both at the word and line levels. This observation is further validated by empirical evidence: when an inversion model trained on Arabic is used to invert embeddings in Meadow Mari (unseen in mT5), the model often generates English, highlighting the influence of pre-training data on embedding inversion attacks. Fig 6 (top) provides a visualization of language confusion at the line level for each language, alongside their corresponding reconstruction performance in BLEU at each step of the evaluation in the crosslingual inversion attacks. It shows directly that lower-resourced languages present lower confusion, especially for non-Latin script languages, whereas they are also more vulnerable in terms of higher reconstruction performance, for example, Gujarati, Punjabi, and Urdu.
@@
Language Confusion Entropy for Train Languages
Our study reveals that inversion performance significantly improves when trained in in-script/in-family settings (ref. Table 13 in the Appendix). Crosslingual inversion performances are comparable to in-script/in-family training when trained in Kazakh (Latin-script) combined with Gujarati and Punjabi, respectively, and language confusion is notably lower. This suggests that while similar languages tend to increase confusion, certain crosslingual combinations can achieve strong performance without the added confusion seen in in-family/in-script training. Overall, these findings highlight the trade-off between inversion performance and language confusion, indicating further optimization is needed to strike the ideal balance between them.
@@ 4.3 Language Confusion and Linguistic Typology
Table 3 shows the best results from KL divergence between language confusion and language similarities based on different language graphs for both LCB and MTEI, using Algorithm 1, the whole results are presented in Table 15 in Appendix.
@@
5 Discussion
Language Confusion Entropy $H_{\mathbf{C}}$ vs. Pass Rate (PR)
As shown in Section 4.1, $H_{\mathbf{C}}$ and PR are correlated but measure distinct aspects of language generation. Importantly, $H_{\mathbf{C}}$ is not intended to replace PR but to offer a deeper quantitative characterization of language confusion.
$H_{\mathbf{C}}$ quantifies behavioral uncertainty through the entropy of a model’s probability distribution across languages. It captures partial confusions (e.g., 40% confidence in the target language, 60% spread across others), even when outputs technically “pass.” PR is a binary performance metric assessing whether outputs meet target language requirements.
Hence, $H_{\mathbf{C}}$ complements PR by revealing hidden confusion patterns: a model with high PR could still exhibit high $H_{\mathbf{C}}$ if probabilities scatter across non-target languages. This makes $H_{\mathbf{C}}$ particularly valuable for diagnosing how models arrive at correct outputs, not just whether they do.
Research Questions